Grade 12th passAlgebra

If m = 2, find the value of:
9 – 5m.................

Harshit Singh
5 Years agoGrade 12th pass

2 Answers

Harshit Singh
5 Years ago
Dear Student

From the question it is given that m = 2
Then, substitute the value of p in the question
= 9–(5 × 2)
= 9–10
= -1

Thanks
Shraddha Singh
5 Years ago
Dear Student,
 
Given : m = 2
Substituting value of m in the equation,
9 - 5m = 9 - 5×2 
= 9-10 = -1
 
Hope it helps!
Thank you ✌️